OpenStudy (anonymous):

A data set consists of the following data points: (2, 4), (4, 7), (5, 12) Please Help!! Will Medal!! The line of best fit has the equation = -1.5 + 2.5x. What does this equation predict for a value of x = 3?

OpenStudy (kirbykirby):

if x = 3, then you plug it in the equation -1.5 + 2.5(3) = -1.5 + 7.5 = 6

OpenStudy (anonymous):

Thank you kindly(: